A) Create an example of a math problem that says to, “Factor fully” and that includes
1
Expert's answer
2020-12-15T14:30:34-0500

Let's use x²-5x+6 = 0 as an example.

To factorise the polynomial, we need two numbers such that when we add them together they give -5 and when we multiply them they give six 6.

Those two numbers are -3 and -2.

x²-3x-2x+6=0

Now, you can factorise.

x(x-3)-2(x-3) =0

The values must be the same in both brackets, otherwise you're wrong.

(x-3)(x-2) =0

x-3 =0 or x-2 =0

Therefore, x =3 or x = 2